OpenStudy (anonymous):

PLEASE HELP --> WILL GIVE MEDAL --> MUST SHOW WORK Task 2 Part 1: Create two radical equations: one that has an extraneous solution, and one that does not have an extraneous solution. Use the equation below as a model. a√x + b + c=d Use a constant in place of each variable a, b, c, and d. You can use positive and negative constants in your equation. __________________________________________________________ Part 2: Show your work in solving the equation. Include the work to check your solution and show that your solution is extraneous. __________________________________________________________ Part 3: Explain why the first equation

OpenStudy (anonymous):

PLEASE HURRY!!! I REALLY NEED HELP!!!

OpenStudy (anonymous):

PLS GUIDE ME STEP BY STEP. I WILL DO WORK!

OpenStudy (anonymous):

A.) a√x+b+c=d has solution: 2√x+5+7=4 has no extraneous solution. 2√x+5+7=11 (b/c 2√x=-1, √x=-.5 x=.25. if you plug it in, you will get a decimal for 2√x)

OpenStudy (anonymous):

Whoops I did it for you without steps :(

OpenStudy (anonymous):

Can you show steps?

OpenStudy (anonymous):

@xo.A

OpenStudy (anonymous):

you got that or no?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

k relax, takes a minute, no more

OpenStudy (anonymous):

ready?

OpenStudy (anonymous):

yes, please! :)

OpenStudy (anonymous):

ok your equation is supposed to look like \[a\sqrt{x+b}+c=d\] right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and you want two equations, one which has a solution and one which does not (which they call "extraneous")

OpenStudy (anonymous):

*thumbs up* yes

OpenStudy (anonymous):

are you still there??? @satellite73

OpenStudy (anonymous):

first we do one that does have a solution, just make up some numbers for \(a,b,c\) and \(d\) for example \[2\sqrt{x+5}+4=10\]

OpenStudy (anonymous):

got hung up,yeah i am here

OpenStudy (anonymous):

*nods head*

OpenStudy (anonymous):

how do i know it will have a solution? answer, because the first step is to subtract \(4\) from both sides, and \(10-4=6\) is positive

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

lets solve it \[2\sqrt{x+5}+4=10\] subtract 4 get \[2\sqrt{x+5}=6\] divide by 2 get \[\sqrt{x+5}=3\] square both sides get \[x+5=9\]

OpenStudy (anonymous):

and last step it so subtract 5 to find \(x=4\)

OpenStudy (anonymous):

that is a solution, one that works, and you can check by substituting in back in to the original equation now want to do one that has only an "extraneous" solution?

OpenStudy (anonymous):

I understand the beginning so far, thx!!! But I'm not so sure about this part, it sounds complicated

OpenStudy (anonymous):

which part ?

OpenStudy (anonymous):

the part with the ''extraneous'' solution, the word is intimidating! :) I'm in a program for 7 graders to do high school math, so ive been a little behind the rest of my class

OpenStudy (anonymous):

ok so lets demystify i t

OpenStudy (anonymous):

it is possible to solve \[\sqrt{x}=-1\]? the answer is "no" why not? because \(\sqrt{x}\) is always greater than or equal to zero, so it cannot be negative not minus 1, minus 2, minus anything

OpenStudy (anonymous):

oh, so does extraneous mean that its impossible, kind of like a null set?

OpenStudy (anonymous):

you can attempt to solve it, say square both sides and get \[x=(-1)^2\\ x=1\] but that is not a solution

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

it is math teacherese for "you got an answer, but it don't work"

OpenStudy (anonymous):

so now we can come up with an example easy just make sure at the first step you will get a negative number on the right of the equal sing

OpenStudy (anonymous):

sign

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

make one up, i will go get a drink and come back in two seconds, see if what you had will give an "extraneous" solution

OpenStudy (anonymous):

\[2\sqrt{x + 4} -10= 7\] Does this make sense

OpenStudy (anonymous):

I meant plus 10

OpenStudy (anonymous):

yes with plus!

OpenStudy (anonymous):

\[2\sqrt{x + 4} +10= 7\]

OpenStudy (anonymous):

yay!!! Thx!!! :)

OpenStudy (anonymous):

first step would be subract \(10\) and get \[2\sqrt{x+4}=-3\]

OpenStudy (anonymous):

which is not possible however let me give you a word of advice

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

part of the problem is actually to "solve" it even though there solution will be extraneous right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so i would start with something like \[2\sqrt{x + 4} +10= 6\] instead of \[2\sqrt{x + 4} +10= 7\] because that way, when you subtrat, you will get an even number

OpenStudy (anonymous):

ok, thx. ill keep that in mind

OpenStudy (anonymous):

then when you divide by 2 as the third step, you won't have an annoying fraction to work with

OpenStudy (anonymous):

second step actually

OpenStudy (anonymous):

you happier now?

OpenStudy (anonymous):

yes!!! :) Ur a great teacher!!!

OpenStudy (anonymous):

(blush) happy to help now i am going to go watch Ma and Pa Kettle on TCM ttytl

OpenStudy (anonymous):

bye, Thx again!!! :)

OpenStudy (anonymous):

yw( again)